One day, the temperature at 2:00 A.M. was 43°F, and by 3:00 P.M. the temperature was 56°F. What was the hourly rate of temperature change?
2°F/h This is the answer I got but it seems wrong can you help?

Answer by funmath(2925)

This one is tricky because of the time change:
Rate=the change in Temperature/change in time
The change in temperature is: 56°F-43°F=13°F
The change in time is the hard part, it might help to think of 2AM as being the 2nd hour of the day and 3PM as being the 15th hour of the day. Then the change in time=15h-2h=13h.
Therefore, the rate of temeperature change is=13°F/13h=1°F/h
Happy Calculating!!!